# Inertial Coordination Games

### v1

@@ A.2  Theorem 1 and Corollary 2
Suppose that $\eta_{t}^{-2}\leq Ct^{p}$ holds for all $t\geq T$. Then we have
$\eta_{t}^{-2}\leq Ct^{p}\implies\eta_{t}^{2}\geq C^{-1}t^{-p}\implies\eta_{t}^{2}+\eta_{t-1}^{2}\geq 2\eta_{t}^{2}\geq 2C^{-1}t^{-p}$
@@
for $t\geq T$. By Lemma 3 (iii), $\mu_{t+1}^{*}-\mu_{t}^{*}$ is decreasing in $t$, which implies
$\mu_{t+1}^{*}-\mu_{t}^{*}\leq t^{-1}\sum_{s=1}^{t}(\mu_{s+1}^{*}-\mu_{s}^{*})=t^{-1}(\mu_{t+1}^{*}-\mu_{1}^{*})\leq t^{-1}.$
@@
Throughout this proof, fix any $q>2$ and (wlog) $\lambda_{0}>1/2$. When convenient, we omit dependencies on these variables. Let “p.m.” denote “on a set of positive (Lebesgue) measure.”
@@
There exists $\underline{C}>0$ such that, for any $T\geq 1$,
@@
Note that $\underline{C}$ does not depend on $t$. We begin with the case $T=1$. Fix any $\Sigma$ with induced posterior precisions $(\eta_{t}^{-2})_{t}$ such that $\eta_{t}^{-2}\geq\underline{C}t^{q}$ for all $t\geq 1$. Then the following implications hold:
@@
$\implies A_{t}\leq\frac{1}{2}\frac{\varphi(\gamma_{1})}{S}t^{-r}\quad\forall t\geq 2$
@@
$\frac{C}{K}=L=\frac{2^{-q}}{2^{-q}+1}\leq\frac{t^{-q}}{t^{-q}+(t-1)^{-q}}\quad\forall\ t\geq 2$
@@
$\text{sign}(L^{\prime}(t))=\text{sign}\Big{(}(t-1)^{-1}-t^{-1}\Big{)}=1.$
@@
$\Phi(\gamma_{t-1}(\epsilon))-\Phi(\gamma_{t}(\epsilon))\geq\varphi(\gamma_{t-1}(\epsilon))(\gamma_{t-1}(\epsilon)-\gamma_{t}(\epsilon))$
$\geq\varphi(\gamma_{1})\Delta_{t}(\epsilon)\geq\frac{\gamma_{t}(\epsilon)}{\gamma_{1}}\varphi(\gamma_{1})\Delta_{t}(\epsilon)\quad\forall t\geq 2$
$\implies\frac{\varphi(\gamma_{1})}{\gamma_{1}}\Delta_{t}(\epsilon)\leq A_{t}(\epsilon)\quad\forall t\geq 2$
because $(\gamma_{t}(\epsilon))_{t}$ is decreasing, and $\gamma_{t}(\epsilon)<\gamma_{1}$ holds for all $t\geq 2$. Also note that
$\lim_{\epsilon\downarrow 0}\frac{\varphi(\gamma_{1})}{\gamma_{1}}\frac{\gamma_{1}-\epsilon}{S}=\frac{\varphi(\gamma_{1})}{S}$
Choose $\epsilon>0$ small enough, independent of $t$, such that
$A_{t}\leq\frac{1}{2}\frac{\varphi(\gamma_{1})}{S}t^{-r}\leq\frac{\varphi(\gamma_{1})}{\gamma_{1}}\frac{\gamma_{1}-\epsilon}{S}t^{-r}\leq A_{t}(\epsilon)\quad\ \forall\ t\geq 2.$
@@
Next, fix any $T\geq 1$, and let $\underline{C}$ be the constant derived above. Fix any $\Sigma$ satisfying $\eta_{t}^{-2}\geq\underline{C}t^{q}$ for all $t\geq T$, and define:
$\tilde{\underline{C}}(\Sigma):=L^{-1}\left[\frac{1}{2}\frac{\varphi(\gamma_{T})}{S}\right]^{-2}\leq\underline{C}$
where the inequality follows because $\gamma_{T}\leq\gamma_{1}$. Hence, $\Sigma$ satisfies
$\eta_{\tilde{t}}^{-2}\geq\underline{C}\{\tilde{t}+(T-1)\}^{q}\geq\underline{C}(\tilde{t})^{q}\geq\tilde{\underline{C}}(\Sigma)(\tilde{t})^{q}\quad\forall\tilde{t}\geq 1$
where $\tilde{t}:=t-(T-1)$. By an exactly analogous argument as the $T=1$ case, $\gamma_{\infty}>0$ holds as desired. ∎
Step 2. To finish the argument, we want to show:
@@
Fix any $C>0$, $T\geq 1$, and $\Sigma$. Fix any $q^{\prime}\in(2,q)$, and choose $T^{\prime}\geq T$ large enough such that
@@
where $\underline{C}(q^{\prime},\lambda_{0})$ is the constant from Lemma 8. By Lemma 8, the following implications hold:
@@ A.3  Proposition 1
Hence, $\gamma_{\infty}\leq\gamma_{\infty}^{\prime}$ which implies
@@
To prove Part (i), assume $\lambda_{0}>(1/2)$ and $\Sigma=\Sigma^{\prime}$. Then,
@@ A.4  Proposition 2
We prove Proposition 2 in two steps. First, we show that for any monotone selection rule with threshold $\mu^{*}\in(c-\lambda_{0},c-(1/2))$, there exists an exponential sequence $(A_{t})_{t}$ with $A_{t}\downarrow 0$ as $t\to\infty$ that implements it, i.e., $\mu_{\infty}^{*}=\mu^{*}$. Second, we show that there exists a noise process $\Sigma=(\sigma_{t})_{t}$ that implements the sequence $(A_{t})_{t}$ constructed in the first step.
@@
Step 1. Fix any $\mu^{*}\in(c-\lambda_{0},c-(1/2))$. Our goal is to construct an exponential sequence $(A_{t})_{t}$ that satisfies $A_{t}\downarrow 0$ as $t\to\infty$, $A_{t+1}^{2}-A_{t+2}^{2}\leq A_{t}^{2}-A_{t+1}^{2}$ for $t\geq 2$, and $\mu_{\infty}^{*}=\mu^{*}$.
@@
where $\gamma_{1}:=\Phi^{-1}(\lambda_{0})$. Hence, it suffices to find an exponential sequence $(A_{t})_{t}$ whose induced $(\gamma_{t})_{t}$ via ($\gamma$-LoM) satisfies $\gamma_{t}\downarrow\gamma^{*}$. By Lemma 7, we can choose $\gamma_{2}\in(\gamma^{*},\gamma_{1})$ close enough to $\gamma^{*}$ such that
@@
By construction, $\gamma_{t}\downarrow\gamma^{*}$ and $A_{t}\downarrow 0$. It remains to show that $(A_{t})_{t}$ is exponential. Note that for all $t\geq 3$,
@@
$\varphi(\gamma_{2})(\gamma_{t-1}-\gamma_{t})\leq\varphi(\gamma_{t-1})(\gamma_{t-1}-\gamma_{t})\leq\Phi(\gamma_{t-1})-\Phi(\gamma_{t})$
$\leq\varphi(\gamma_{t})(\gamma_{t-1}-\gamma_{t})\leq\varphi(\gamma^{*})(\gamma_{t-1}-\gamma_{t}),$
where the second and the third inequalities follow from the mean value theorem. Hence,
$\underbrace{4\frac{\varphi(\gamma_{2})}{\gamma_{2}}(\gamma_{2}-\gamma^{*})}_{=:c}2^{-t}\leq\frac{\Phi(\gamma_{t-1})-\Phi(\gamma_{t})}{\gamma_{2}}\leq A_{t}$
$\leq\frac{\Phi(\gamma_{t-1})-\Phi(\gamma_{t})}{\gamma^{*}}\leq\underbrace{4\frac{\varphi(\gamma^{*})}{\gamma^{*}}(\gamma_{2}-\gamma^{*})}_{=:C}2^{-t}$
as desired, where the second and the third inequalities follow from $\gamma^{*}<\gamma_{t}\leq\gamma_{2}$ for all $t\geq 2$.
Finally, we want to show that the sequence $(A_{t})_{t}$ constructed above satisfies $A_{t+1}^{2}-A_{t+2}^{2}\leq A_{t}^{2}-A_{t+1}^{2}$ for $t\geq 2$ if we take $\gamma_{2}$ sufficiently close to $\gamma_{2}$. We defined the sequence of $A_{t}$ so that $A_{2}\geq\sqrt{2}A_{3}$ holds. Hence, we have
@@ A.5  Proposition 3
First, we show that $T(\sigma)\to\infty$ as $\sigma\downarrow 0$. Notice that $T(\sigma)$ is weakly decreasing in $\sigma$ because for any $\sigma^{\prime}>\sigma$, $\theta\geq\mu^{*}_{T(\sigma^{\prime})}(\sigma^{\prime})>\mu^{*}_{T(\sigma^{\prime})}(\sigma)$ implies $T(\sigma)\geq T(\sigma^{\prime})$ by the definition of $T(\sigma)$. Hence, $T(\sigma)$ either converges to some constant or diverges to $\infty$ as $\sigma\downarrow 0$. Suppose for a contradiction that $\lim_{\sigma\downarrow 0}T(\sigma)=T$ holds for some $T>0$. Since $T(\sigma)\uparrow T$ as $\sigma\downarrow 0$, for any $\sigma$, we have
@@
Since the lower bound on $\theta-\mu_{t}^{*}$ holds uniformly for $t\in[0,\alpha T(\sigma)]$, for any $\epsilon>0$, there exists $\underline{\sigma}_{\alpha}>0$ such that for all $\sigma<\underline{\sigma}_{\alpha}$ and all $t\in[0,\alpha T(\sigma)]$, $\lambda_{t}\geq 1-\epsilon$.
@@
$\leq\frac{1}{\sqrt{2\pi}\sigma}\left|(\sqrt{t+1}-\sqrt{t})\theta-\sqrt{t+1}\mu_{t+1}^{*}+\sqrt{t}\mu_{t}^{*}\right|$
@@
$<\frac{\sqrt{t+1}-\sqrt{t}}{\sqrt{2\pi}\sigma}|\theta|+\frac{\sqrt{t+1}}{\sqrt{2\pi}\sigma t}+\frac{c(\sqrt{t+1}-\sqrt{t})}{\sqrt{2\pi}\sigma}$
$\leq\frac{1}{\sqrt{2\pi}\sigma}\left\{(\sqrt{2}-1)(|\theta|+c)+\sqrt{2}\right\},$

### v2

@@ A.2  Theorem 1 and Corollary 2
First, note that for any $\epsilon>2$, $\lim_{t\to\infty}t^{2-\epsilon}=0$. Hence, $\eta_{t}^{-2}=O(t^{2-\epsilon})$ implies $\lim_{t\to\infty}\eta_{t}^{-2}=0$. By our assumption of limit learning, the implication we want to show is vacuously true.
Hence, fix any $\epsilon\in(0,2)$, and let $p=2-\epsilon$. Suppose there exist $C,T$ such that $\eta_{t}^{-2}\leq Ct^{p}$ holds for all $t\geq T$. Then we have
$\eta_{t}^{-2}\leq Ct^{p}\implies\eta_{t}^{2}\geq C^{-1}t^{-p}\implies\eta_{t}^{2}+\eta_{t-1}^{2}\geq 2\eta_{t}^{2}\geq 2C^{-1}t^{-p}$ ($\eta_{t}^{2}\leq\eta_{t-1}^{2}$)
@@
for all $t\geq T$. By Lemma 3 (iii), $|\mu_{t+1}^{*}-\mu_{t}^{*}|=\mu_{t+1}^{*}-\mu_{t}^{*}$ is decreasing in $t$, which implies
$\mu_{t+1}^{*}-\mu_{t}^{*}\leq t^{-1}\sum_{s=1}^{t}(\mu_{s+1}^{*}-\mu_{s}^{*})=t^{-1}(\mu_{t+1}^{*}-\mu_{1}^{*})\leq t^{-1}.$ ($\mu_{t+1}^{*}-\mu_{1}^{*}\leq 1$)
@@
Throughout this proof, we denote $q=2+\epsilon$ and when convenient, we omit dependencies on the variables $q>2$ and (wlog) $\lambda_{0}>1/2$. Let “p.m.” denote “on an interval of positive (Lebesgue) measure.”
@@
For any $q>2$ and $\lambda_{0}>1/2$, there exists $\underline{C}>0$ (depending on $q$ and $\lambda_{0}$) such that, for any $T\geq 1$,
@@
Note that $\underline{C}$ does not depend on $t$, but does depend on $q,\lambda_{0}$. We begin with the case $T=1$. Fix any $\Sigma$ with induced posterior precisions $(\eta_{t}^{-2})_{t}$ such that $\eta_{t}^{-2}\geq\underline{C}t^{q}$ for all $t\geq 1$. Then the following implications hold:
@@
$\implies A_{t}\leq\frac{1}{2}\frac{\varphi(\gamma_{1})}{S}t^{-r}\quad\forall t\geq 2$ (1)
@@
$\frac{C}{K}=L=\frac{2^{-q}}{2^{-q}+1}\leq\frac{t^{-q}}{t^{-q}+(t-1)^{-q}}\quad\forall\ t\geq 2.$
@@
$\frac{d}{dt}\left(\frac{t^{-q}}{t^{-q}+(t-1)^{-q}}\right)=q(t(t-1))^{-q}\Big{(}(t-1)^{-1}-t^{-1}\Big{)}\geq 0$
@@
$\Phi(\gamma_{t-1}(\epsilon))-\Phi(\gamma_{t}(\epsilon))$ $\geq\varphi(\gamma_{t-1}(\epsilon))(\gamma_{t-1}(\epsilon)-\gamma_{t}(\epsilon))$ ($0<\gamma_{t}(\epsilon)<\gamma_{t-1}(\epsilon)$ $\forall t\geq 2$)
$\geq\varphi(\gamma_{1})\Delta_{t}(\epsilon)$ ($0<\gamma_{t-1}(\epsilon)\leq\gamma_{1}$ and $\gamma_{t-1}(\epsilon)-\gamma_{t}(\epsilon)=\Delta_{t}(\epsilon)$ $\forall t\geq 2$)
$\geq\frac{\gamma_{t}(\epsilon)}{\gamma_{1}}\varphi(\gamma_{1})\Delta_{t}(\epsilon)\quad\forall t\geq 2$ ($\frac{\gamma_{t}(\epsilon)}{\gamma_{1}}\in(0,1)$ $\forall t\geq 2$)
$\implies\frac{\varphi(\gamma_{1})}{\gamma_{1}}$ $\Delta_{t}(\epsilon)\leq A_{t}(\epsilon)\quad\forall t\geq 2$ (2)
where we also use the fact that $\varphi(x)$ is decreasing for $x\geq 0$. Also note that
$\lim_{\epsilon\downarrow 0}\frac{\varphi(\gamma_{1})}{\gamma_{1}}\frac{\gamma_{1}-\epsilon}{S}=\frac{\varphi(\gamma_{1})}{S}>\frac{1}{2}\frac{\varphi(\gamma_{1})}{S}$
Hence, we can choose $\epsilon>0$ small enough, independent of $t$, such that
[…]
@@
Next, fix any $T\geq 1$, and let $\underline{C}$ be the constant derived above (which does not depend on $T$). Fix any $\Sigma$ satisfying $\eta_{t}^{-2}\geq\underline{C}t^{q}$ for all $t\geq T$, and define:
$\tilde{\underline{C}}(\Sigma):=L^{-1}\left[\frac{1}{2}\frac{\varphi(\gamma_{T})}{S}\right]^{-2}\leq\underline{C}.$
where we note that $0\leq\gamma_{T}\leq\gamma_{1}$ implies $\varphi(\gamma_{T})\geq\varphi(\gamma_{1})$. Hence, if we define $\tilde{t}:=t-(T-1)$ and $\tilde{\eta}_{\tilde{t}}^{-2}=\eta_{t}^{-2}$, $\Sigma$ satisfies
$\tilde{\eta}_{\tilde{t}}^{-2}\geq\underline{C}\{\tilde{t}+(T-1)\}^{q}\geq\underline{C}(\tilde{t})^{q}\geq\tilde{\underline{C}}(\Sigma)(\tilde{t})^{q}\quad\forall\tilde{t}\geq 1$
By an exactly analogous argument as the $T=1$ case, $\gamma_{\infty}>0$ holds as desired. ∎
Step 2. To finish the argument, we want to show: for any $q>2$ and (wlog) $\lambda_{0}>1/2$,
@@
Fix any $q>2$, $C>0$, $T\geq 1$, and $\Sigma$. Fix any $q^{\prime}\in(2,q)$, and choose $T^{\prime}\geq T$ large enough such that
@@
where $\underline{C}(q^{\prime},\lambda_{0})$ is the constant from Lemma 8. Then, the following implications hold:
@@ A.3  Proposition 1
Hence, $\gamma_{\infty}\leq\gamma_{\infty}^{\prime}$ holds, which implies
@@
To prove Part (i), assume $\lambda_{0}>(1/2)$ and $\Sigma=\Sigma^{\prime}$. Then, by the same inductive argument as above, we have
@@ A.4  Proposition 2
We prove Proposition 2 in two steps. First, we show that for any threshold $\mu^{*}\in(c-\lambda_{0},c-(1/2))$, there exists a sequence $(A_{t})_{t}$ with $A_{t}\downarrow 0$ as $t\to\infty$ that implements it, i.e., $\mu_{\infty}^{*}=\mu^{*}$. Second, we show that there exists a noise process $\Sigma=(\sigma_{t})_{t}$ that implements the sequence $(A_{t})_{t}$ constructed in the first step.
@@
Step 1. Fix any $\mu^{*}\in(c-\lambda_{0},c-(1/2))$. Our goal is to construct a sequence $(A_{t})_{t}$ that satisfies $A_{t}\downarrow 0$ as $t\to\infty$, $A_{t+1}^{2}-A_{t+2}^{2}\leq A_{t}^{2}-A_{t+1}^{2}$ for $t\geq 2$, and $\mu_{\infty}^{*}=\mu^{*}$.
@@
where $\gamma_{1}:=\Phi^{-1}(\lambda_{0})$. Hence, it suffices to find a sequence $(A_{t})_{t}$ whose induced $(\gamma_{t})_{t}$ via ($\gamma$-LoM) satisfies $\gamma_{t}\downarrow\gamma^{*}$. By Lemma 6 and Lemma 7, we can choose $\gamma_{2}\in(\gamma^{*},\gamma_{1})$ close enough to $\gamma^{*}$ such that
@@
By construction, $\gamma_{t}\downarrow\gamma^{*}$ and $A_{t}\downarrow 0$. It remains to show that $(A_{t})_{t}$ converges exponentially, which is a fact we use in the next step. Note that for all $t\geq 3$,
@@
$\varphi(\gamma_{2})(\gamma_{t-1}-\gamma_{t})$
$\leq\varphi(\gamma_{t-1})(\gamma_{t-1}-\gamma_{t})\leq$ $\Phi(\gamma_{t-1})-\Phi(\gamma_{t})\leq\varphi(\gamma_{t})(\gamma_{t-1}-\gamma_{t})$ (Mean value theorem)
$\leq$ $\varphi(\gamma^{*})(\gamma_{t-1}-\gamma_{t})$
Hence,
$\underbrace{4\frac{\varphi(\gamma_{2})}{\gamma_{2}}(\gamma_{2}-\gamma^{*})}_{=:c}2^{-t}\leq\frac{\Phi(\gamma_{t-1})-\Phi(\gamma_{t})}{\gamma_{2}}\leq A_{t}$ ($\gamma_{t}\leq\gamma_{2}$)
$\leq\frac{\Phi(\gamma_{t-1})-\Phi(\gamma_{t})}{\gamma^{*}}\leq\underbrace{4\frac{\varphi(\gamma^{*})}{\gamma^{*}}(\gamma_{2}-\gamma^{*})}_{=:C}2^{-t}$ ($\gamma^{*}<\gamma_{t}$)
as desired.
Finally, we want to show that the sequence $(A_{t})_{t}$ constructed above satisfies $A_{t+1}^{2}-A_{t+2}^{2}\leq A_{t}^{2}-A_{t+1}^{2}$ for $t\geq 2$ if we take $\gamma_{2}$ sufficiently close to $\gamma^{*}$. We defined the sequence of $A_{t}$ so that $A_{2}\geq\sqrt{2}A_{3}$ holds. Hence, we have
@@ A.5  Proposition 3
First, we show that $T(\sigma)\to\infty$ as $\sigma\downarrow 0$. Notice that $T(\sigma)$ is weakly decreasing in $\sigma$ because, by the proof of Proposition 1(ii), for any $\sigma^{\prime}>\sigma$, $\theta\geq\mu^{*}_{T(\sigma^{\prime})}(\sigma^{\prime})>\mu^{*}_{T(\sigma^{\prime})}(\sigma)$. Hence, $T(\sigma)\geq T(\sigma^{\prime})$ by the definition of $T(\sigma)$. Hence, $T(\sigma)$ either converges to some constant or diverges to $\infty$ as $\sigma\downarrow 0$. Suppose for a contradiction that $\lim_{\sigma\downarrow 0}T(\sigma)=T$ holds for some $T>0$. Since $T(\sigma)\uparrow T$ as $\sigma\downarrow 0$, for any $\sigma$, we have
@@
Since our lower bound on $\lambda_{t}$ holds uniformly for $t\in[0,\alpha T(\sigma)]$, for any $\epsilon>0$, there exists $\underline{\sigma}_{\alpha}>0$ such that for all $\sigma<\underline{\sigma}_{\alpha}$ and all $t\in[0,\alpha T(\sigma)]$, $\lambda_{t}\geq 1-\epsilon$.
@@
$\leq\frac{1}{\sqrt{2\pi}\sigma}\left|(\sqrt{t+1}-\sqrt{t})\theta-\sqrt{t+1}\mu_{t+1}^{*}+\sqrt{t}\mu_{t}^{*}\right|$ ($\Phi^{\prime}(\cdot)\leq 1/\sqrt{2\pi}$)
@@
$<\frac{\sqrt{t+1}-\sqrt{t}}{\sqrt{2\pi}\sigma}|\theta|+\frac{\sqrt{t+1}}{\sqrt{2\pi}\sigma t}+\frac{c(\sqrt{t+1}-\sqrt{t})}{\sqrt{2\pi}\sigma}$ ($\mu_{t}^{*}<c$ and $\mu_{t+1}^{*}-\mu_{t}^{*}<t^{-1}$)
$\leq\frac{1}{\sqrt{2\pi}\sigma}\left\{(\sqrt{2}-1)(|\theta|+c)+\sqrt{2}\right\}.$